OpenStudy (anonymous):

A cyclist travels at a distance of 120km from Town A to Town B at an average speed of x km/h. If he increases his speed by 1 km/h, he would have saved 30 minutes. Find x.

OpenStudy (perl):

distance = (average rate) * (time elapsed) 120 = x * t 120 = (x+1) (t - .5) solve this system

OpenStudy (jjuden):

120/x+1 = 120/x - 1/2h 120/x+1 = 240-x/2x 240x = (240-x)(x+1) 240x = 240x + 240 - x^2 -x x+x^2 = 240 15+15^2 = 240 therefore x= 15

OpenStudy (jjuden):

orr x= (-1 + sqrt(1^2 - 4(1)(-240)))/(2*1) x= (-1 + sqrt(961)) / 2 x= (-1 + 31) /2 x= 30/2 = 15

OpenStudy (perl):

do you have any questions?

OpenStudy (perl):

distance = (average rate) * (time elapsed) 120 = x * t 120 = (x+1) (t - 1/2) to solve this system, solve for t in the first equation t = 120/x then 120 = (x+1)( 120/x - 1/2) then distribute 120 = x*120/x + 120/x - x/2 - 1/2 120 = 120 + 120/x - x/2 - 1/2 0 = 120/x - x/2 - 1/2 multiply both sides by 2x 0 = 240 - x^2 - x 0 = (x - 15)(x+16) 0 = 240 - x^2 - 2x